I need a refresher in order to work with linear equations, please help! The problem I am trying to solve is:

x=y+3
Y=3-x

Could this be the answer? 5= 2+3 and 2= 3-1

Found 2 solutions by usyim88hk, cleomenius:
Answer by usyim88hk(157)

Yes, you are right but you can solve it with a algebratical way

Answer by cleomenius(959)

x = y + 3
y = 3 - x
x - y = 3
x + y = 3
Add the 2 equations.
2x = 6
x = 3
==============================
Now plug into the first equation to solve for y.
3 = 0 + 3.
y = 0
check in the second equation.
0 = 3 - 3.
This does check,
Cleomenius.